OpenStudy (anonymous):

[Properties of exponents] --Simplify. Your answer should contain only positive exponents. 1) 4v^3 [times] vu^2 ----and---- 2) 4a^3b^2 [times] 3a^-4b^-3. Thanks in advance c:

OpenStudy (anonymous):

Do you know the rule for multiplying two things with the same base? \[x^2*x^4\] for example?

OpenStudy (anonymous):

2x^6 :s?

OpenStudy (anonymous):

Actually, just x^6

OpenStudy (anonymous):

Oops.

OpenStudy (anonymous):

The rule is simple to add the exponents. Why? Because \[x^2 * x^4 = (x*x)*(x*x*x*x) = x*x*x*x*x*x = x^6\]

OpenStudy (anonymous):

Now if there are different bases, we can't use that rule. So something like: \[x^3 * y^2\] we wouldn't do anything to. We'd just leave it like that.

OpenStudy (anonymous):

So that's what we're going to do with these. Look at the first one and let me see you try it.

OpenStudy (anonymous):

Use the rule for things that have the same base, but leave other things alone.

OpenStudy (anonymous):

Sorry I fell asleep, lol.

OpenStudy (anonymous):

Uhmm, 4v^3-1u^2 = 4v^4u^2 c:

OpenStudy (anonymous):

Good. 4v^4u^2 is correct

OpenStudy (anonymous):

Thanks so much c: